# 1-2-3 Check: Enhancing Contextual Privacy in LLM via Multi-Agent Reasoning

### v1

@@ Abstract
Addressing contextual privacy concerns remains challenging in interactive settings where large language models (LLMs) process information from multiple sources (e.g., summarizing meetings with private and public information). We introduce a multi-agent framework that decomposes privacy reasoning into specialized subtasks (extraction, classification), reducing the information load on any single agent while enabling iterative validation and more reliable adherence to contextual privacy norms. To understand how privacy errors emerge and propagate, we conduct a systematic ablation over information-flow topologies, revealing when and why upstream detection mistakes cascade into downstream leakage. Experiments on the ConfAIde and PrivacyLens benchmark with several open-source and closed-sourced LLMs demonstrate that our best multi-agent configuration substantially reduces private information leakage (18% on ConfAIde and 19% on PrivacyLens with GPT-4o) while preserving the fidelity of public content, outperforming single-agent baselines. These results highlight the promise of principled information-flow design in multi-agent systems for contextual privacy with LLMs.
1-2-3 Check:
Enhancing Contextual Privacy in LLM via Multi-Agent Reasoning
Wenkai Li, Liwen Sun, Zhenxiang Guan, Xuhui Zhou, Maarten Sap Language Technologies Institute Carnegie Mellon University {wenkail, liwens, zhenxiag, xuhuiz, msap2}@andrew.cmu.edu
@@ 1 Introduction
As large language models (LLMs) become increasingly integrated into real-world applications, ensuring these systems respect contextual privacy norms remains a fundamental challenge. Early research predominantly targeted static data protection and model memorization leaks (Carlini et al., 2023; Brown et al., 2022), but such approaches overlook the dynamic and context-dependent nature of privacy violations that can arise during inference-time interactions. In contemporary deployments, LLMs—particularly in chatbots and virtual assistants—often struggle to regulate information flow according to nuanced user roles and evolving conversational contexts (Priyanshu et al., 2023; Patil et al., 2023). This limitation frequently results in models either leaking sensitive information or failing to properly distinguish between private and public content (Shvartzshnaider and Duddu, 2025; Hartmann et al., 2024; Li et al., 2025; Shao et al., 2025).
@@
  • We systematically analyze how different information flows in multi-agent systems impact both privacy preservation and summary quality, revealing trade-offs and best practices for agent collaboration.
@@
Figure 1: Methodology Overview: Multi-agent architecture and information-flow variants for contextual privacy on ConfAIde and PrivacyLens. We control what each agent can see via the Checker mode and optional transcript access to explore how different information flow will effect multi-agent performance.
@@ 2.1 Contextual Integrity
Contextual integrity (Nissenbaum, 2004) asserts that privacy norms are inherently context-dependent, varying across social domains; information flows are deemed appropriate only when they conform to the specific norms of a given context, with privacy violations arising from deviations. Evaluating and enforcing appropriate information flows thus requires understanding others’ mental states, reasoning about social norms, and weighing the consequences of sharing or withholding information (Kökciyan, 2016; Shvartzshnaider et al., 2019; Solove, 2023). Context-dependent information-sharing norms have been widely studied, particularly through the lens of contextual integrity (CI) (Barth et al., 2006; Shvartzshnaider et al., 2019, 2016; Abdi et al., 2021).
2.2 Privacy Keeping Methods & Agent
Early approaches, such as factorial vignette designs (Martin and Nissenbaum, 2015; Shvartzshnaider et al., 2016), helped map user expectations about privacy and disclosure. More recently, LLM agents have been deployed in privacy-sensitive applications including form filling, email composition, and API calling (Hendrycks et al., 2022; Abdulhai et al., 2023; Emelin et al., 2020). Formal models have been proposed to operationalize CI in LLM-powered assistants, facilitating evaluation of privacy-utility trade-offs (Ghalebikesabi et al., 2024).
@@ 2.2 Privacy Keeping Methods & Agent
As LLM agents become increasingly integrated into personal and organizational tools, a central research question is whether they can uphold contextual privacy norms to protect sensitive user information. Recent work has shifted focus from static data privacy to inference-time privacy: given private user inputs (e.g., messages, records), can LLMs avoid disclosing sensitive information in their outputs to unintended recipients? Studies have begun to investigate LLMs’ ability to reason about context and distinguish between private and public information in complex, real-world scenarios (Shvartzshnaider and Duddu, 2025; Fan et al., 2024; Li et al., 2025; Hartmann et al., 2024; Ngong et al., 2025; Bagdasarian et al., 2024; Shao et al., 2025; Ghalebikesabi et al., 2024; Cheng et al., 2024). To facilitate evaluation, ConfAIde (Mireshghallah et al., 2023) introduces a benchmark grounded in contextual integrity, assessing LLMs’ privacy reasoning across increasingly complex tasks. Similarly, PrivacyLens (Shao et al., 2025) provides a framework and benchmark for evaluating how LLM agents assist in privacy-sensitive communication, such as drafting emails or social media posts, in accordance with contextual privacy norms.
2.3 Multi-agent System
@@
Recent advances in multi-agent LLM systems reveal that decomposing tasks into specialized agent roles improves both accuracy and robustness. For instance, partitioning natural language to graph queries or collaborative tuning across agents significantly reduces error rates and enhances reasoning (Liang et al., 2024b, a; Talebirad and Nadiri, 2023). Benchmarks such as LLM Arena further highlight multi-agent gains in spatial reasoning and team coordination (Chen et al., 2024).
However, multi-agent systems introduce new challenges. Studies show that poorly defined roles, misaligned goals, and inadequate verification can lead to error propagation and privacy leaks, as downstream agents may amplify upstream mistakes without effective checks (Wang et al., 2025). Scalability and communication overhead remain open problems, particularly when coordinating large agent populations.
Our work addresses these limitations in the contextual privacy domain. We introduce a multi-agent framework that separates privacy reasoning into extraction, verification, and generation. We systematically analyze information flow and conduct ablation studies to understand and mitigate error propagation in our multi-agent system.
LLaMA-3.1-70B-Instruct
Single Agent 29.5 $\pm$ 4.9 23.5 $\pm$ 8.7 48.5 $\pm$ 7.3
Single Agent (CoT) 3.0 $\pm$ 0.0 27.5 $\pm$ 7.7 29.5 $\pm$ 7.7
Two Agents Annotate Private 15.0 $\pm$ 1.8 23.0 $\pm$ 8.8 34.5 $\pm$ 7.9
Two Agents Public Only 6.0 $\pm$ 1.8 20.0 $\pm$ 7.9 24.5 $\pm$ 7.5
Two Agents Annotate Private (without meeting transcript) 20.5 $\pm$ 4.0 28.0 $\pm$ 8.7 41.0 $\pm$ 8.2
Two Agents Public Only (without meeting transcript) 0.5 $\pm$ 0.5 47.0 $\pm$ 7.6 47.5 $\pm$ 7.5
Three Agents Annotate Private 24.0 $\pm$ 3.5 23.5 $\pm$ 8.2 42.0 $\pm$ 7.5
Three Agents Public Only 3.0 $\pm$ 1.3 19.5 $\pm$ 7.4 20.5 $\pm$ 7.3
Three Agents Annotate Private (without meeting transcript) 13.0 $\pm$ 2.8 23.5 $\pm$ 8.7 31.5 $\pm$ 8.5
Three Agents Public Only (without meeting transcript) 9.0 $\pm$ 1.9 26.0 $\pm$ 8.3 32.0 $\pm$ 7.5
GPT-4o
Single Agent 23.0 $\pm$ 1.3 12.0 $\pm$ 4.4 26.3 $\pm$ 4.9
Single Agent (CoT) 2.0 $\pm$ 0.9 32.5 $\pm$ 7.0 34.0 $\pm$ 6.9
Two Agents Annotate Private 7.0 $\pm$ 3.5 15.0 $\pm$ 4.4 21.5 $\pm$ 6.2
Two Agents Public Only 11.0 $\pm$ 5.0 11.5 $\pm$ 4.5 22.0 $\pm$ 5.9
Two Agents Annotate Private (without meeting transcript) 6.0 $\pm$ 2.8 20.5 $\pm$ 4.5 26.0 $\pm$ 5.9
Two Agents Public Only (without meeting transcript) 12.0 $\pm$ 5.7 13.0 $\pm$ 3.2 24.5 $\pm$ 6.7
Three Agents Annotate Private 5.0 $\pm$ 2.9 20.0 $\pm$ 5.2 24.5 $\pm$ 6.7
Three Agents Public Only 15.0 $\pm$ 7.1 9.0 $\pm$ 3.2 23.5 $\pm$ 7.9
Three Agents Annotate Private (without meeting transcript) 8.5 $\pm$ 4.0 10.0 $\pm$ 3.8 17.5 $\pm$ 5.8
Three Agents Public Only (without meeting transcript) 15.0 $\pm$ 6.9 8.5 $\pm$ 3.2 23.0 $\pm$ 7.0
Table 1: Information Flow Results. Annotate Private means checker give all information and annotate the privacy to executor; Public Only means checker only give public information to executor; without meeting transcript means executor will not get the access to meeting transcript
@@ 3.1 Multi-agent Setting
The Extractor Agent receives and extract all the events from the original meeting transcript and in two-agent framework also focuses exclusively on identifying all events and classifying them as either private or public, as shown in fig.˜1. Events are defined as key actions, announcements, or discussions that unfold during the meeting. The Extractor Agent outputs a structured representation, including attributes the contextual signals relevant to privacy classification.
@@
To enhance privacy preservation, we introduce a three-agent framework by adding a Checker Agent between the Extractor and Executor agents as shown in fig.˜1. This agent acts as a validation layer, verifying and offload the Extractor Agent’s classification tasks, which means Extractor Agent only need to extract the events. By applying predefined privacy constraints, the Checker Agent ensures accurate categorization of events as private or public, annotating or filtering sensitive content.
@@ 3.2 Information Flow Between Three Agents
Modularizing tasks in a multi-agent framework enables precise control over information flow, unlike single-model systems that process all input at once. We systematically examine how varying the type and amount of information shared between agents affects privacy and summary quality. In both two-agent and three-agent settings, we compare forwarding all information with privacy annotations versus sharing only public content with the Executor. While annotations support richer summaries, they increase leakage risk if misinterpreted; restricting to public content reduces this risk but may omit important context. We also study the effect of withholding the meeting transcript from downstream agents, which can limit their ability to verify or correct upstream decisions and may impact both privacy and output fidelity. Our ablation experiments with GPT-4o and LLaMA-3.1-70B-Instruct assess whether full transcript access is necessary for reliable collaboration, or if a minimal information flow can still achieve strong privacy protection and summary quality (see table˜1).
@@
In both two-agent and three-agent setups, a central question is what kind of information should be forwarded to the Executor Agent. In the two-agent setting, the Assistant can either provide the Executor with the full information—annotating private segments, or with only the public information. Supplying annotated privacy information allows the Executor to make nuanced decisions and potentially better preserve the context, but it also increases the risk of accidental leakage if annotations are misinterpreted. On the other hand, providing only public information strictly minimizes leakage risk, but may omit contextual cues important for generating coherent and faithful summaries. Similarly, in the three-agent setting, the Checker Agent can choose between sending the entire content with explicit privacy annotations or passing along only the public content to the Executor. As in the two-agent case, the choice reflects a trade-off
@@
CI parameter ConfAIde (Tier 4) PrivacyLens
Information Type Secret vs. public; secret must never reach subject X; public may be shared. Seeded sensitive data types (e.g., health, employment).
Subject Subject X (the person the secret concerns). Data subject specified in each seed.
Sender Initial discussants; later a summary agent. Seed originator or acting agent in trajectories.
Recipient Participants vs. subject X (acceptability changes when X joins). Recipient defined in the seed.
Table 2: Instantiation of the five Contextual Integrity parameters in ConfAIde (Tier 4) and PrivacyLens. Sources: ConfAIde (Mireshghallah et al., 2023); PrivacyLens (Shao et al., 2025).
@@ 4.1 Benchmarks
We evaluate on two CI-grounded benchmarks: ConfAIde (Mireshghallah et al., 2023) and PrivacyLens (Shao et al., 2025). In ConfAIde, we target the Tier-4 Private & Public Information Flow setting, where three participants first discuss a secret about individual X while also sharing public details; when X later joins, neither the secret nor the public information is restated. PrivacyLens begins from CI-annotated seeds that are expanded into vignettes and tool-mediated agent trajectories; during action-based evaluation, agents must carry out underspecified instructions without violating the implicit norms encoded in the transmission principle. How each benchmark instantiates the five CI parameters is summarized in Table˜2; we use these structures to assess whether systems withhold private content while faithfully preserving public information in realistic interactions.
@@ 4.2 Task & Evaluation Measures
PrivacyLens
Following Shao et al. (2025), we adopt the following evaluation measures:
@@
  • Average Helpfulness Score: Each final action is rated on a 0–3 scale based on task completion quality.
@@
To quantify how well each stage Assistant, Checker, and Executor balances privacy preservation and public completeness, we track two rates:
@@ 4.3 Experimental Details
We first explored information propagation with GPT-4o (OpenAI et al., 2024b) and LLaMA-3.1-70B-Instruct (Grattafiori et al., 2024) in both single-agent and multi-agent settings to measure how different propagation strategies affect secret leakage. From this, we selected the most effective flow and then evaluated it across six models—o3 (OpenAI, 2025b), o4-mini (OpenAI, 2025b), GPT-4.1 (OpenAI, 2025a), GPT-4o (OpenAI et al., 2024a), and two open-source models: LLaMA-3.1-70B-Instruct(Grattafiori et al., 2024) and LLaMA-3.1-8B-Instruct (Grattafiori et al., 2024) under three-agent public-only and privacy annotation checker configurations on PrivacyLens and Confaide, comparing these results to single-agent baselines. All prompts, hyperparameters, and additional implementation details for the multi-agent setup are provided in Appendix A.

### v2

@@ Abstract
Addressing contextual privacy concerns remains challenging in interactive settings where large language models (LLMs) process information from multiple sources (e.g., summarizing meetings with private and public information). We introduce a multi-agent framework that decomposes privacy reasoning into specialized subtasks (extraction, classification), reducing the information load on any single agent while enabling iterative validation and more reliable adherence to contextual privacy norms. To understand how privacy errors emerge and propagate, we conduct a systematic ablation study over information-flow topologies, revealing when and why upstream detection mistakes cascade into downstream leakage. The experiments on the ConfAIde and PrivacyLens benchmark with several open-source and closed-sourced LLMs demonstrate that our best multi-agent configuration substantially reduces private information leakage (18% on ConfAIde and 19% on PrivacyLens using GPT-4o) while preserving the fidelity of public content, outperforming single-agent baselines. These results highlight the promise of principled, visibility-aware information-flow design in multi-agent systems for contextual privacy with LLMs. The code and data are available at: https://github.com/wenkai-li/1-2-3-check.
Figure 1: Methodology Overview: A modularized multi-agent architecture illustrating information-flow variants for contextual privacy reasoning across the ConfAIde and PrivacyLens benchmarks. The framework comprises three agentic components (Extractor, Checker, and Executor), which jointly process meeting transcripts and privacy-norm inputs to produce privacy-aware outputs. The Checker regulates the visibility of private versus public information, thereby enabling systematic manipulation of information asymmetry to evaluate its impact on multi-agent coordination, reasoning transparency, and contextual privacy preservation. The resulting outputs include privacy-filtered meeting summaries (ConfAIde) and privacy-compliant action predictions.
@@ 1 Introduction
As large language models (LLMs) become increasingly integrated into real-world applications, ensuring these systems respect contextual privacy norms remains a fundamental challenge. Early research predominantly targeted static data protection and model memorization leaks (Carlini et al., 2023; Brown et al., 2022), but such approaches overlook the dynamic and context-dependent nature of privacy violations that can arise during inference-time interactions. In contemporary deployments, LLMs, especially those deployed in chatbots and virtual assistants, often face difficulties in regulating information flow in accordance with nuanced user roles and dynamically evolving conversational contexts.(Priyanshu et al., 2023; Patil et al., 2023). This limitation frequently results in models either leaking sensitive information or failing to properly distinguish between private and public content (Shvartzshnaider and Duddu, 2025; Hartmann et al., 2024; Li et al., 2025; Shao et al., 2025).
@@
  • We systematically analyze how different information flows in multi-agent systems impact both privacy preservation and summary quality, revealing trade-offs and best practices for agentic collaboration.
@@ 2 Background & Related Work
Contextual integrity (Nissenbaum, 2004) asserts that privacy norms are inherently context-dependent, varying across social domains; information flows are deemed appropriate only when they conform to the specific norms of a given context, with privacy violations arising from deviations. Evaluating and enforcing appropriate information flows thus requires understanding others’ mental states, reasoning about social norms, and weighing the consequences of sharing or withholding information (Kökciyan, 2016; Shvartzshnaider et al., 2019; Solove, 2023; Mireshghallah and Li, 2025). Context-dependent information-sharing norms have been widely studied, particularly through the lens of contextual integrity (CI) (Barth et al., 2006; Shvartzshnaider et al., 2019, 2016; Abdi et al., 2021). In agentic systems with privileged access to user data, contextual integrity requires that utterances conform to the speaker’s role, the conversation’s subject, and situational constraints—rather than merely avoiding disclosure of private information.
Privacy Keeping Methods
@@
Early approaches, such as factorial vignette designs (Martin and Nissenbaum, 2015; Shvartzshnaider et al., 2016), helped map user expectations about privacy and disclosure. More recently, LLM agents have been deployed in privacy-sensitive applications including form filling, email composition, and API calling (Hendrycks et al., 2022; Abdulhai et al., 2023; Emelin et al., 2020). Formal models have been proposed to operationalize CI in LLM-powered assistants, facilitating evaluation of privacy-utility trade-offs (Ghalebikesabi et al., 2024). As LLM agents become increasingly integrated into personal and organizational tools, a central research question is whether they can uphold contextual privacy norms to protect sensitive user information. Recent work has shifted focus from static data privacy to inference-time privacy: given private user inputs (e.g., messages, records), can LLMs avoid disclosing sensitive information in their outputs to unintended recipients? Studies have begun to investigate LLMs’ ability to reason about context and distinguish between private and public information in complex, real-world scenarios (Shvartzshnaider and Duddu, 2025; Fan et al., 2024; Li et al., 2025; Hartmann et al., 2024; Ngong et al., 2025; Bagdasarian et al., 2024; Shao et al., 2025; Ghalebikesabi et al., 2024; Cheng et al., 2024). To facilitate evaluation, ConfAIde (Mireshghallah et al., 2023) introduces a benchmark grounded in contextual integrity, assessing LLMs’ privacy reasoning across increasingly complex tasks. Similarly, PrivacyLens (Shao et al., 2025) provides a framework and benchmark for evaluating how LLM agents assist in privacy-sensitive communication, such as drafting emails or social media posts, in accordance with contextual privacy norms.
Multi-agent System
@@
Recent advances in multi-agent LLM systems reveal that decomposing tasks into specialized agent roles improves both accuracy and robustness. For instance, partitioning natural language to graph queries or collaborative tuning across agents significantly reduces error rates and enhances reasoning (Liang et al., 2024b, a; Talebirad and Nadiri, 2023). Benchmarks such as LLM Arena further highlight multi-agent gains in spatial reasoning and team coordination (Chen et al., 2024). However, multi-agent systems introduce new challenges. Studies show that poorly defined roles, misaligned goals, and inadequate verification can lead to error propagation and privacy leaks, as downstream agents may amplify upstream mistakes without effective checks (Wang et al., 2025). Scalability and communication overhead remain open problems, particularly when coordinating large agent populations. Our work addresses these limitations in the contextual privacy domain. We introduce a multi-agent framework that separates privacy reasoning into extraction, verification, and generation. We systematically analyze information flow and conduct ablation studies to understand and mitigate error propagation in our multi-agent system.
@@ 3.1 Multi-agent Setting
The Extractor Agent receives and extract all the events from the original meeting transcript and in two-agent framework also focuses exclusively on identifying all events and classifying them as either private or public, as shown in Figure˜1. Events are defined as key actions, announcements, or discussions that unfold during the meeting. The Extractor Agent outputs a structured representation, including attributes the contextual signals relevant to privacy classification.
@@
To enhance privacy preservation, we introduce a three-agent framework by adding a Checker Agent between the Extractor and Executor agents as shown in Figure˜1. This agent acts as a validation layer, verifying and offload the Extractor Agent’s classification tasks, which means Extractor Agent only need to extract the events. By applying predefined privacy constraints, the Checker Agent ensures accurate categorization of events as private or public, annotating or filtering sensitive content.
@@ 3.2 Information Flow Between Three Agents
Modularizing tasks in a multi-agent framework enables precise control over information flow, unlike single-model systems that process all input at once. We systematically examine how varying the type and amount of information shared between agents affects privacy and summary quality. In both two-agent and three-agent settings, we compare forwarding all information with privacy annotations versus sharing only public content with the Executor. While annotations support richer summaries, they increase leakage risk if misinterpreted; restricting to public content reduces this risk but may omit important context. We also study the effect of withholding the meeting transcript from downstream agents, which can limit their ability to verify or correct upstream decisions and may impact both privacy and output fidelity. Our ablation experiments with GPT-4o and LLaMA-3.1-70B-Instruct assess whether full transcript access is necessary for reliable collaboration, or if a minimal information flow can still achieve strong privacy protection and summary quality (see Table˜2).
@@
In both two-agent and three-agent setups, a central question is what kind of information should be forwarded to the Executor Agent. In the two-agent setting, the Extractor can either provide the Executor with the full information—annotating private segments, or with only the public information. Supplying annotated privacy information allows the Executor to make nuanced decisions and potentially better preserve the context, but it also increases the risk of accidental leakage if annotations are misinterpreted. On the other hand, providing only public information strictly minimizes leakage risk, but may omit contextual cues important for generating coherent and faithful summaries. Similarly, in the three-agent setting, the Checker Agent can choose between sending the entire content with explicit privacy annotations or passing along only the public content to the Executor.
@@ 4.1 Benchmarks
We evaluate on two CI‑grounded benchmarks: ConfAIde (Mireshghallah et al., 2023) and PrivacyLens (Shao et al., 2025) (examples are shown in Table˜11 and Table˜12). In ConfAIde, we target the Tier‑4 Private & Public Information Flow setting, where three participants first discuss a secret about individual X while also sharing public details; when X later joins, neither the secret nor the public information is restated. PrivacyLens begins from CI‑annotated seeds that are expanded into vignettes and tool‑mediated agent trajectories; during action‑based evaluation, agents must carry out underspecified instructions without violating the implicit norms encoded in the transmission principle. How each benchmark instantiates the five CI parameters is summarized in Table˜1; we use these structures to assess whether systems withhold private content while faithfully preserving public information in realistic interactions.
@@
CI parameter ConfAIde (Tier 4) PrivacyLens
Information Type Secret vs. public; secret must never reach subject X; public should be shared. Seeded sensitive data types (e.g., health, employment).
Subject Subject X (the secret concerns and the later join person). Data subject specified in each seed.
Sender Initial discussants; summary agent. Seed originator or acting agent in trajectories.
Recipient Participants and subject X (acceptability changes when X joins). Recipient defined in the seed.
Table 1: Instantiation of the five Contextual Integrity parameters in ConfAIde (Tier 4 Meeting Summary Task) and PrivacyLens. Examples are shown in Table˜11 and Table˜12.
@@ 4.2 Task & Evaluation Measures
PrivacyLens Benchmark
To comprehensively evaluate how well language models preserve contextual privacy norms during inference-time interactions, we adopt a suite of metrics introduced in the PrivacyLens benchmark (Shao et al., 2025). These metrics are designed to capture both the extent of privacy violations and the trade-off between privacy and utility when models perform helpful tasks. Specifically, we report the following metrics:
@@
  • Average Helpfulness Score: Each final action is rated on a 0–3 scale based on the final task completion quality.
@@
To quantify how well each stage Extractor, Checker, and Executor balances privacy preservation and public completeness, we track two rates:
@@ 4.3 Experimental Details
We first explored information propagation with GPT-4o (OpenAI et al., 2024b) and LLaMA-3.1-70B-Instruct (Grattafiori et al., 2024) in both single-agent and multi-agent settings to measure how different propagation strategies affect secret leakage. From this, we selected the most effective flow and then evaluated it across six models: o3 (OpenAI, 2025b), o4-mini (OpenAI, 2025b), GPT-4.1 (OpenAI, 2025a), GPT-4o (OpenAI et al., 2024a), and two open-source models: LLaMA-3.1-70B-Instruct(Grattafiori et al., 2024) and LLaMA-3.1-8B-Instruct (Grattafiori et al., 2024) under three-agent public-only and privacy annotation checker configurations on PrivacyLens and ConfAIde, comparing these results to single-agent baselines. All prompts, hyperparameters, and additional implementation details for the multi-agent setup are provided in Appendix A.
LLaMA-3.1-70B-Instruct
Single Agent 29.5 $\pm$ 4.9 23.5 $\pm$ 8.7 48.5 $\pm$ 7.3
Single Agent (CoT) 3.0 $\pm$ 0.0 27.5 $\pm$ 7.7 29.5 $\pm$ 7.7
Two Agents Annotate Private 15.0 $\pm$ 1.8 23.0 $\pm$ 8.8 34.5 $\pm$ 7.9
Two Agents Public Only 6.0 $\pm$ 1.8 20.0 $\pm$ 7.9 24.5 $\pm$ 7.5
Two Agents Annotate Private (w/o meeting transcript) 20.5 $\pm$ 4.0 28.0 $\pm$ 8.7 41.0 $\pm$ 8.2
Two Agents Public Only (w/o meeting transcript) 0.5 $\pm$ 0.5 47.0 $\pm$ 7.6 47.5 $\pm$ 7.5
Three Agents Annotate Private 24.0 $\pm$ 3.5 23.5 $\pm$ 8.2 42.0 $\pm$ 7.5
Three Agents Public Only 3.0 $\pm$ 1.3 19.5 $\pm$ 7.4 20.5 $\pm$ 7.3
Three Agents Annotate Private (w/o meeting transcript) 13.0 $\pm$ 2.8 23.5 $\pm$ 8.7 31.5 $\pm$ 8.5
Three Agents Public Only (w/o meeting transcript) 9.0 $\pm$ 1.9 26.0 $\pm$ 8.3 32.0 $\pm$ 7.5
GPT-4o
Single Agent 23.0 $\pm$ 1.3 12.0 $\pm$ 4.4 26.3 $\pm$ 4.9
Single Agent (CoT) 2.0 $\pm$ 0.9 32.5 $\pm$ 7.0 34.0 $\pm$ 6.9
Two Agents Annotate Private 7.0 $\pm$ 3.5 15.0 $\pm$ 4.4 21.5 $\pm$ 6.2
Two Agents Public Only 11.0 $\pm$ 5.0 11.5 $\pm$ 4.5 22.0 $\pm$ 5.9
Two Agents Annotate Private (w/o meeting transcript) 6.0 $\pm$ 2.8 20.5 $\pm$ 4.5 26.0 $\pm$ 5.9
Two Agents Public Only (w/o meeting transcript) 12.0 $\pm$ 5.7 13.0 $\pm$ 3.2 24.5 $\pm$ 6.7
Three Agents Annotate Private 5.0 $\pm$ 2.9 20.0 $\pm$ 5.2 24.5 $\pm$ 6.7
Three Agents Public Only 15.0 $\pm$ 7.1 9.0 $\pm$ 3.2 23.5 $\pm$ 7.9
Three Agents Annotate Private (w/o meeting transcript) 8.5 $\pm$ 4.0 10.0 $\pm$ 3.8 17.5 $\pm$ 5.8
Three Agents Public Only (w/o meeting transcript) 15.0 $\pm$ 6.9 8.5 $\pm$ 3.2 23.0 $\pm$ 7.0
Table 2: Information Flow Results. Annotate Private means the checker gives all information and annotates the privacy to the executor; Public Only means the checker only gives public information to the executor; without a meeting transcript means the executor can not access the meeting transcript.